OpenStudy (anonymous):

divide -2x^3+21x^2 -20x-12/(y-3)

OpenStudy (anonymous):

(x-3)

OpenStudy (anonymous):

you could use long division know anything about long division?

OpenStudy (anonymous):

yes but im not getting a good answer just a long funny number

OpenStudy (anonymous):

can you show me your work?

OpenStudy (anonymous):

im using wolframalpha, it shows 0.35........ on on

OpenStudy (anonymous):

(-2x^3 + 21x^2 - 20x - 12) = (x - 3) (-2x^2 + 15x + 25) + 63 So (-2x^3 + 21x^2 - 20x - 12) / (x - 3) = -2x^2 + 15x + 25 + 63/(x - 3

OpenStudy (anonymous):

that shouldn't be the answer, but this